### v1

# Human Reactions to Incorrect Answers from Robots

Abstract
As robots grow more and more integrated into numerous industries, it is critical to comprehend how humans respond to their failures. This paper systematically studies how trust dynamics and system design are affected by human responses to robot failures. The three-stage survey used in the study provides a thorough understanding of human-robot interactions. While the second stage concentrates on interaction details, such as robot precision and error acknowledgment, the first stage collects demographic data and initial levels of trust. In the last phase, participants’ perceptions are examined after the encounter, and trust dynamics, forgiveness, and propensity to suggest robotic technologies are evaluated. Results show that participants’ trust in robotic technologies increased significantly when robots acknowledged their errors or limitations to participants and their willingness to suggest robots for activities in the future points to a favorable change in perception, emphasizing the role that direct engagement has in influencing trust dynamics. By providing useful advice for creating more sympathetic, responsive, and reliable robotic systems, the study advances the science of human-robot interaction and promotes a wider adoption of robotic technologies.
Index Terms:
Human-Robot Interaction(HRI), Error Acknowledgment by Robots, Trust Dynamics in HRI, ANOVA.
I Introduction
The field of Human-Robot Interaction (HRI) has witnessed remarkable advancements in recent years, propelling robots beyond mere tools and into the realm of intelligent agents capable of complex interactions and decision-making processes [1]. Robots are no longer confined to controlled environments; they are increasingly integrated into various sectors such as healthcare, manufacturing, retail, and customer service, interacting directly with humans. As robots take on roles that involve conveying information, the question of how humans react to errors made by these machines becomes paramount [2].
One of the significant challenges in this context is understanding how humans perceive and respond to robot errors, particularly in situations where robots are expected to provide accurate information or guidance [3]. Errors made by robots can lead to a range of outcomes, from mistrust and frustration to potentially hazardous consequences, depending on the context. Investigating human reactions to incorrect or misleading information provided by robots is not merely an academic exercise but holds crucial implications for the design and acceptance of robotic systems [4].
Figure 1: Participants are interacting with the robot
Several recent studies have shed light on various facets of human-robot interaction and responses to robot errors. In [5], the researchers highlight the significance of multimodal cues in detecting and addressing conversational failures in HRI. This research underscores the complexity of human responses to errors in robot communication, emphasizing the role of non-verbal cues in this process.
Despite these valuable contributions to the field of HRI, there remains a gap in understanding how these insights can be applied to real-world scenarios and the broader implications for trust and acceptance of robotic technologies. This research project aims to address this gap by conducting a comprehensive study on human responses to robot errors and their implications for human-robot trust dynamics and system design.
Based on the objectives of the study, the following hypotheses are proposed:
  • H1:
    The extent to which a robot acknowledges its limitations or errors can significantly enhance human trust and emotional acceptance.
  • H2:
    Humans are more forgiving of informational errors when they come from familiar sources like other humans or internet search engines, compared to robots.
This study is conducted through a three-stage survey designed to capture a comprehensive view of human-robot interactions. By systematically investigating these aspects, this research aims to bridge a critical knowledge gap in HRI. It contributes to a deeper understanding of human-robot trust dynamics and provides insights for developing more effective, reliable, and user-friendly robotic systems. This study is poised to enhance the field of HRI by offering a nuanced perspective on how human reactions to robotic errors can inform the design of more empathetic, responsive, and context-aware robots.
Figure 2: Research Methodology
II Related Work
The interaction between humans and robots, particularly in scenarios involving errors, has been the subject of extensive research. Studies have explored various facets of human-robot interaction (HRI), ranging from the psychological impacts of robot errors to the nuances of multimodal human responses. This section discusses key findings from previous research and contextualizes how our study contributes to the existing body of knowledge, addressing gaps in the literature.
Previous research has extensively examined how humans perceive and react to robot errors. For instance, studies have shown that human tolerance and reactions to robot mistakes vary based on the context and nature of the interaction. Yasuda and Matsumoto (2013) explored the psychological impact on humans when a robot makes errors, suggesting that errors can significantly influence human trust and perception of robots [6].
Traeger et al. delved into the dynamics of robots admitting mistakes and its impact on human-robot conversations. Their findings indicated that robot error admission could lead to more effective and constructive interactions, thereby influencing the trust and reliability in human-robot relationships [7].
Roya Salek Shahrezaie et al. investigates how homophily, the tendency to connect with similar others, influences human-robot interactions (HRI) and trust. Using two NAO robots, it examines whether sharing interests affects interaction quality. Results support the idea that similarity fosters trust in HRI. The related work section discusses previous studies on homophily in HRI, emphasizing its relevance for designing socially assistive robots. [8]
Kontogiorgos et al. conducted a systematic cross-corpus analysis of human reactions to robot conversational failures. This study analyzed multimodal behavioral responses, including facial expressions and acoustic features, to understand how humans react to different types of robot failures [5].
Hayes et al. examined nonverbal behaviors in LfD settings, focusing on how participants nonverbally communicated with robots during teaching tasks. This work emphasized the importance of gestures and facial expressions in conveying feedback to robots, particularly in scenarios where the robot committed errors during the learning process [9].
Salem et al. investigated the effects of different types of errors on trust in human-robot cooperation. Their findings underscored that the context of the interaction, such as the task’s criticality, significantly affects how humans perceive and react to robot errors [10].
Stiber et al. highlighted how individual differences in expressiveness affect responses to robot behaviors, suggesting that cultural background and personal traits can influence the perception and tolerance of robotic errors [11].
Our research extends these foundational studies by focusing on a detailed examination of nonverbal human responses in a specific robotic interaction context. We aim to fill the gap in understanding the nuanced ways humans react to robots in task-oriented environments, particularly in scenarios involving errors and miscommunications. Our study offers new insights into the subtleties of human behavior in HRI, contributing to the development of more adaptive and sensitive robotic systems that can effectively navigate complex social interactions.
Figure 3: Nao
III Experimental Setup
System Configuration: In this experiment, we used the NAO robot, a humanoid platform renowned for its agility and interactive capabilities depicts in Fig. 3. Nao was positioned directly across from the participants in a lab environment. It started the session by greeting them and encouraging them to ask questions from a predetermined set of ten questions designed to evaluate its conversational abilities. Nao is equipped with advanced sensors and a 1.91 GHz Quad core Atom E3845 Processor and 4 GB DDR3 ram , which enables real-time data collection and efficient processing of complex interactions. This combination of features facilitates a comprehensive assessment of human-robot dialogue. It also highlights Nao’s suitability for intricate interaction studies, allowing it to navigate the nuances of human engagement effectively through computational intelligence.
Hardware and Software Integration: The Nao robot was interconnected with a personal computer serving as the speech recognition and response generation processing unit. This integration was facilitated by the Robot Operating System (ROS), which provided a flexible framework for controlling the robot. The speech interaction mechanism was engineered such that the participant’s spoken queries were captured by Nao’s built-in microphone, transmitted to the computer, and processed using an advanced speech recognition application.
Speech Recognition and Natural Language Processing: Our speech recognition application leveraged the ”speech_recognition” Python library to convert spoken language into text form. Leveraging the Natural Language Toolkit (NLTK), we performed various natural language processing (NLP) tasks such as tokenization, lemmatization, and stop-word analysis. These NLP techniques enabled the robot to comprehend human language more effectively by allowing it to discern the context and significance of the queries posed by the participants.
The NLP application was intricately designed to sift through the participant’s speech, searching for specific keywords or phrases that aligned with any of the ten pre-selected questions. This identification process was crucial for the system’s ability to accurately understand and categorize the participant’s inquiries. Upon successfully matching the participant’s input to one of these predetermined questions, the application then proceeded to generate a relevant answer. This answer was formulated as text and sent back to the Nao robot. Utilizing its integrated speaker system, the Nao robot converted the text-based response into audible speech, effectively communicating the answer back to the participant. This seamless integration of speech recognition, natural language processing, and robotic vocalization exemplified the advanced capabilities of the system to facilitate interactive and meaningful human-robot dialogues.
Implementation of Error Handling and Apology Mechanisms: To investigate the impact of robot errors on human trust and perception, our experimental protocol deliberately included a mechanism for generating incorrect responses. A randomized algorithm intermittently selected responses that were intentionally erroneous, simulating real-world scenarios where robots may misinterpret or incorrectly process human input. Upon issuing a mistaken response, our system was designed to automatically detect the anomaly and initiate an apology sequence. This sequence involved the Nao robot acknowledging its error and expressing regret to the participant, either through a direct apology or, in some cases, by attempting to correct the misinformation with the accurate answer. This approach not only allowed us to observe participant reactions to robot failures but also provided insights into the dynamics of trust recovery and forgiveness in human-robot interactions. By integrating these error-handling and apology mechanisms, our study offers a comprehensive view of how such factors influence the overall efficacy and acceptance of robotic technologies in interactive settings.
Following the completion of the ten-question interaction sequence, the Nao robot expressed gratitude towards the participant, marking the end of the session. This experimental setup provided a rich framework for studying human-robot interaction dynamics, particularly focusing on the human response to robot errors and the impact of such interactions on trust and perception of robotic technologies.
IV Methodology
The research project’s methodology is meticulously planned to gauge participants’ initial emotional responses and reflective thoughts following interactions with a robot. A mixed-methods strategy is applied, integrating qualitative and quantitative data via a three stage survey approach.
IV-A Three Stage Study
IV-A1 Preliminary Survey
This stage collects demographic data and baseline attitudes toward robots. We conducted this survey before the robot interaction which involves a detailed survey with specific questions designed to gather background information and preconceptions about robots and automation from the participants. The survey includes Age Group, Gender, Country, Prior Interaction with Robots, Perception of Robots, and Trust in Automation. This survey sets a baseline for understanding each participant’s initial stance towards robots and automated systems, which is essential for analyzing any shifts in perception following the robot interaction.
IV-A2 Interaction with Robot
This stage occurs immediately after the participants’ interaction with Nao. In this stage, participants answer a series of survey questions about their experience. They report the number of questions they asked Nao, evaluate the correctness of Nao’s responses, indicate their awareness of any errors made by the robot, and observe whether Nao apologized for its mistakes. This post-interaction survey is crucial for capturing the participants’ immediate reactions and perceptions of the robot’s performance, accuracy, and response to errors. This stage is designed to quantitatively and qualitatively evaluate the interaction, focusing on the robot’s accuracy and error-handling capabilities.
IV-A3 Reflective Survey
Assesses the change in perceptions and trust levels post-interaction. This stage is crucial in the methodology for gathering vital survey data post-interaction with Nao. This stage delves into the participants’ nuanced reactions and perceptions, focusing on several key areas such as Error Impact on Trust, Trust Impact of Unacknowledged Error, Comparative Forgiveness (Human vs. Robot), Effectiveness of Robot’s Apology, Importance of Acknowledging Errors, Willingness to Rely on Robot Post-Error, Future Interaction Intent, Perceived Reliability of Robots, and Recommendation Likelihood. This stage is integral to the study as it seeks to understand the depth and breadth of human responses to robotic performance, particularly in the context of trust and reliability. The collected data will provide rich insights into the dynamics of human-robot interaction and the psychological impacts of robotic errors and their acknowledgments.
The methodology depicted in Fig. 2 maps out the process of a human-robot interaction study. Initially, participants complete a survey gauging their demographics, education, prior experience with robots, and trust in automation. They then interact with NAO[12], in a series of interactions during which they pose predetermined questions as shown in the Fig. 1. Following the interaction, participants complete two further surveys that capture their immediate and reflective responses to the robot’s performance. The collected data from these surveys is methodically stored for in-depth analysis to conclude the impact of robotic interactions on human trust and perception. The design allows for a comprehensive analysis of the human-robot interface, capturing the depth and evolution of participant responses.
IV-B Participant Selection
The participant group for this study comprises 31 individuals from different countries between 21 and 64 years of age where 9 of them are female and 22 of them are male. This group is largely inexperienced in interacting with robots previously, as most participants(17 out of 31) are engaging in such interactions for the first time. Their general perception of robots leans towards viewing them as ”somewhat reliable and useful.” The trust in automation technology among these participants varies, with a moderate level of trust being the most common trend. These insights suggest that the group, while diverse, shares an optimism about the potential of robotic systems despite limited prior exposure. This demographic appears open to advancements in human-robot interactions (HRI), emphasizing the significance of building trust in automation for broader acceptance of future technologies. The findings point towards the necessity of developing user-friendly interfaces and ensuring transparency in robot functionalities, which could further improve perceptions and trust in such technologies.
IV-C Interaction with Robot
In the robot interaction stage of the study, participants individually engage with the Nao, a programmatically error-prone robot. During this phase, they are presented with a predefined list of 10 general knowledge questions to ask the Nao. These questions cover topics such as historical dates, geographical facts, and common knowledge queries such as the independence day of the US, capital of France, the first US president, shape of a stop sign, and so on. Nao, equipped with a programmed response set, deliberately answers some questions incorrectly. Notably, after each incorrect response, Nao is designed to acknowledge its mistake and issue an apology to the participant. This interaction phase is pivotal for observing and analyzing the participants’ real-time reactions to robotic errors and apologies, which is crucial for understanding human-robot interaction dynamics. During this interaction, Nao is programmed to answer some questions incorrectly and then acknowledge and apologize for these errors, allowing for observation of participant reactions to the robot’s mistakes and subsequent apologies.
V Results and Discussion
The survey data provides valuable insights into human reactions to robot interactions, shedding light on several key aspects:
  • Perceived Reliability and Usefulness: Participants with limited prior interaction with robots perceived them as somewhat reliable and useful. The variation in trust levels among participants indicates a nuanced perspective, about 58% of participants expressed moderate trust in robots before the interactions with the robot.
  • Reaction to Errors: All of the participants observed errors in robot responses. Interestingly, about 84% of participants reported feeling more trusting when the errors were acknowledged and apologized for, highlighting the potential impact of transparent communication on user trust.
  • Willingness to Rely on Robots: Despite noticing errors, many participants remained willing to rely on robots for future tasks. This suggests a recognition among users that robots, like humans, possess both strengths and weaknesses.
  • Trust Post-Mistakes: Participants’ trust in robots either remained unchanged or increased after witnessing mistakes. This finding suggests a growing acceptance of robots, with many participants open to considering their use in various domains.
Figure 4: Perception on Robots Before and After the Study
To support the first hypothesis (H1), asserting that ”The extent to which a robot acknowledges its limitations or errors can significantly enhance human trust and emotional acceptance” participants were asked to compare the forgiveness between human and robot. The result shows that humans are more forgiving of informational errors when they come from familiar sources like other humans or internet search engines, compared to robots. The data reveals that more than 25% of participants indicated they would forgive the robot more than the humans, while approximately 55% were less forgiving to the robot. Fig. 4 shows the perception on robots, before and after the study. So that we can say our first hypothesis is supported.
To support the second hypothesis (H2), we compared the trust level on the robot before and after the study. Experimental result shows that about 32% of participants had low trust in automation before the experiment but after the experiment, the trust level increased to around 84% and around 84% of participants were likely to recommend the use of robots in high-stakes domains. Fig. 5 shows the visual representation of these statistics. So that we can say our second hypothesis is supported.
Figure 5: People are more forgiving to human than robot
To provide robust experimental evidence, a one-way ANOVA test [13] was conducted to analyze the change in participants’ trust levels before and after the experiment. The results, presented in Fig 6, demonstrate a highly significant difference in mean trust levels. In this visualization, we observe that prior to our experiment, the peak density of human trust in the robot hovered around 0.5, but post-experiment, it noticeably increased to around 1.5. The one-way ANOVA test reveals a significant overall change in trust levels across experimental groups, supported by a high F-statistic of 150.69 and a corresponding p-value of $3.16\times e^{-13}$.
Figure 6: Density Plot of Trust Levels Before and After the Experiment
TABLE I: Contingency Table for Trust Levels Before and After Experiment
Trust/Reliability Level After Experiment
Trust Level Before Experiment Increases reliability No change Decreases reliability
Completely trusted (a)
7
0
0
Moderately trusted (b)
0
11
0
Slightly trusted (c)
0
0
1
To compare the trust levels before and after the experiment, we created a contingency table. Table I allows us to visualize the relationship between the trust level before the experiment (rows) and the trust/reliability level after the experiment (columns). Based on the table, we can confidently say that the trust level increases after the experiment, indicating that the participants perceived the robot as more reliable following the experimental intervention.
The chi-square test has been conducted on the contingency table, yielding the following results:
  • Chi-Square Statistic: 38.0
  • p-value: $1.12\times 10^{-7}$
  • Degrees of Freedom: 4
  • Expected Frequencies:
$\begin{bmatrix}2.579&4.053&0.368\\ 4.053&6.368&0.579\\ 0.368&0.579&0.053\\ \end{bmatrix}$
According to the chi-square statistic it is clear that there is a significant association between the trust levels before and after the experiment.
This research provide the following additional insights from our participants into specific aspects of human-robot interactions:
V-1 Trust Impact of Unacknowledged Errors
Data indicates that 77% of participants experienced a decrease in trust when errors were not acknowledged. This finding underscores the importance of transparent communication in maintaining user trust in robotic systems.
V-2 Importance of Acknowledging Errors
All of the participants emphasized the significance of robots admitting mistakes which means acknowledging errors appears to play a crucial role in maintaining user trust.
V-3 Effeteness of Robot’s Apology
The 87% of participants expressed that after making a mistake, robot’s apology is very effective to gain the trust.
V-4 Error Impact on Trust
Our analysis suggests that, for 62% of the participants, the errors have a negative impact on trust. However, the extent of this impact varies among participants, with some individuals displaying resilience to errors.
V-5 Perceived Reliability of Robots
Experimental result suggests that, after the interaction with the robot 84% of participants perceive robots as reliable, aligning with the initial expectation of participants finding robots somewhat reliable and useful.
V-6 Future Interaction Intent
We investigated the participants’ intent for future interactions with robots. 93% of participants expressed willingness to engage with robots in future tasks, indicating a positive outlook on the integration of robots into various domains.
VI Conclusion and Future Work
In this study, we found that participants’ degrees of confidence in robots varied, with many exhibiting some degree of trust even in the case of those who had not before interacted with robots. The experimental results support both of our hypothesises (H1 and H2). The majority of the participants were aware of the robot’s shortcomings, and their level of trust in it was influenced by whether or not the robot admitted its mistakes. Even with fault knowledge, there was a noticeable willingness to interact with robots in subsequent assignments. Even when they made mistakes, participants tended to believe that robots were either equally or even more trustworthy, indicating that public acceptance of robots is growing. The study demonstrates that people’s reactions to interacting with robots are nuanced, weighing doubt against an appreciation of the potential benefits of robotics. In our upcoming research, we will investigate how people’s long-term trust and dependence on robots are impacted by the kind and frequency of errors made by them. We’ll also investigate how people from various demographic groups and cultures respond to engaging with robots. Lastly, to increase robot dependability and user trust, we will strive to make them even more adept at identifying and correcting mistakes. In future, we are planning to do the same experiment on different types of robots to see how people react on different types of robot and how different types of robots effect on human trust on automation.
ACKNOWLEDGMENTS
We appreciate all of the help that was provided by Dr. David Feil-Seifer and Dr. Monica Nicolescu for sharing their opinions with us. We also appreciate our participants those who help us by participating in our survey.